lauren paddles her kayak 8 miles upstream in 4 hours, she turns around and paddles downstream to her starting point in 2 hours. what is the rate at which lauren paddles in still water? what is the rate of the rivers current?
Answer by TimothyLamb(4379)

---
s = d / t
---
x = kayak speed
y = water speed
---
upstream:
x - y = 8/4 = 2
---
downstream:
x + y = 8/2 = 4
---
x - y = 2
x + y = 4
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = kayak speed = 3 mph
y = water speed = 1 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php